OpenStudy (anonymous):

-2x+5y=19 3x-4=y Solving System of Equation using Substitution.

OpenStudy (tyteen4a03):

We know that y = 3x-4, so just substitute this fact into the other equation, then solve for x. Use the value of x to solve y.

OpenStudy (anonymous):

-2x+ 5(3x-4)=19 -2x+ 15x-20=19 13x-20= 19 cann you find x now?

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

thanx

OpenStudy (whpalmer4):

Sometimes you won't have a set of equations where one of them is quite so convenient, and you'll have to do a little more algebra to get one variable in terms of another. If your second equation had been 3x-4 = 3y, for example, you wouldn't have quite as obvious a choice of what to substitute where, but the approach would be the same. Pick one variable, solve an equation for it in terms of the other(s), plug it into the other equation(s), repeat as needed. Then go backward to work out the other variables.